Show all work: Dalia just paid a dividend of $1. The rate of return is 8%. The dividend is expected to grow at a rate of 3% for the next two years. After the first two years, the dividend will grow at a constant rate of 4%. What is the value of this stock today?
Answer #1

Given:

Rate of return = 8%

Dividend (D0) = $1

Growth rate = 3%

Constant growth = 4%

Dividend in year 1 (D1) = 1 ×1.03 = 1.03

Dividend in year 2 (D2) = 1.03 × 1.03 = 1.0609

Dividend after year 2 (D3) ,= 1.0609 × 1.04 = 1.1033

Price of the stock = D3 / rate of return - growth rate

= 1.1033 / 0.08 - 0.04

= 1.1033 / 0.04

= 27.5834

Value of the stock today = D1 / (1+r) + D2 / (1+r)^2 + price of stock / (1+r)^2

= 1.03 / (1+0.08) + 1.0609/(1+0.08)^3 + 27.5834/(1+0.08)^2

= 1.03/1.08 + 1.0609/(1.08)^2 + 27.5834/(1.08)^2

= 0.9537 + 1.0609 / 1.1664 + 27.5834 / 1.1664

= 0.9537 + 0.9096 + 23.6483

= $ 25.51